find the length of time for $25000 to triple when it is invested at 8% annual interest rate, compounded continuously.
Answer by stanbon(75887)

find the length of time for $25000 to triple when it is invested at 8% annual interest rate, compounded continuously.
---
A(t) = P*e^(rt)
----
Any amount (P) will triple when e^(0.08t) = 3
---------
Take the natural log to get:
0.08t = ln(3)
t = [ln(3)]/0.08
---
t = 13.73.. years
===========================================
Cheers,
Stan H.
===========